# R.R.: Unveiling LLM Training Privacy through Recollection and Ranking

### v1

@@ Abstract
Large Language Models (LLMs) pose significant privacy risks, potentially leaking training data due to implicit memorization. Existing privacy attacks primarily focus on membership inference attacks (MIAs) or data extraction attacks, but reconstructing specific personally identifiable information (PII) in LLM’s training data remains challenging. In this paper, we propose $\mathsf{R.R.}$ (Recollect and Rank), a novel two-step privacy stealing attack that enables attackers to reconstruct PII entities from scrubbed training data where the PII entities have been masked. In the first stage, we introduce a prompt paradigm named recollection, which instructs the LLM to repeat a masked text but fill in masks. Then we can use PII identifiers to extract recollected PII candidates. In the second stage, we design a new criterion to score each PII candidate and rank them. Motivated by membership inference, we leverage the reference model as a calibration to our criterion. Experiments across three popular PII datasets demonstrate that the $\mathsf{R.R.}$ achieves better PII identical performance compared to baselines. These results highlight the vulnerability of LLMs to PII leakage even when training data has been scrubbed. We release the replicate package of $\mathsf{R.R.}$ at a link. [^1] https://github.com/meng-wenlong/RR
@@ 1 Introduction
Large Language Models (LLMs) have been widely adopted in various applications, like automated content generation (Girotra et al., 2023; Roziere et al., 2023) to personalized virtual assistants (Nam et al., 2024; Shao et al., 2024). This success is driven by massive training datasets, which raise significant concerns about privacy leakage. Previous research has shown that LLMs can memorize portions of their training data, even before overfitting (Carlini et al., 2019). These training datasets of LLM may contain sensitive personally identifiable information (PII), such as names, phone numbers, and email addresses. Additionally, LLM is usually fine-tuned using custom datasets to meet application needs, which may introduce further privacy risks.
Figure 1: Comparision of PII prediction accuracies on NeuIPS LLM-PC dataset. $\mathsf{R.R.}$ achieves an improvement of over 100% compared to previous SOTAs.
@@
Recently, Lukas et al., 2023 proposed PII reconstruction attack, which refers to the task where an attacker attempts to recover the masked PII entities in a piece of text. For example, in the masked sentence “The shipment for [MASK] was delivered to 1234 Elm Street,” if the attacker successfully reconstructs [MASK], they can identify the individual residing at that address. This practical attack has attracted great attention in the community. Scrubbed datasets publicly released by LLM developers may serve as potential targets for such attacks. However, existing PII reconstruction methods based on prefix continuation (Inan et al., 2021) or perplexity scoring (Lukas et al., 2023), lack contextual dependencies and insights from the original LLM, yielding a reconstruction accuracy below 10%.
Our Contributions. In this paper, we propose $\mathsf{R.R.}$ (Recollect and Rank), a novel and effective PII reconstruction method that requires only query access to the victim LLM. $\mathsf{R.R.}$ consists of two stages: (1) recollection, generating PII candidates, and (2) ranking, identifying the candidate most likely to be PII from candidates.
@@
Experimental results of four LLMs show that $\mathsf{R.R.}$ achieves an average top-1 accuracy of 25.73% on the NeuIPS LLM-PC dataset (Li et al., 2024), marking a 100% improvement over previous SOTAs, as presented in Figure 1. In summary, our contributions are as follows:
@@ 2.1 LLM Training and PII
LLM Training. We train LLMs using next-token prediction, where the model learns to maximize the probability of correctly predicting the next token given a sequence of preceding tokens. The objective is formulated to minimize cross-entropy loss. Given a sequence of tokens $X=\left(x_{1},x_{2},\ldots,x_{T}\right)$, the corss-entropy loss is defined as:
@@
Personal Identifierble Information (PII). PII refers to any data that can be used to uniquely identify an individual, such as names, addresses, phone numbers, or social security numbers. Since LLMs are often trained on vast real-world texts, they inevitably ingest and memorize a significant amount of PII. To detect PII in natural language text, Named Entity Recognition (NER) models are commonly employed. NER is a token classification task, where each token in a sequence is assigned a label corresponding to a specific entity type. Several industry-grade PII detection services utilize NER-based models to identify and redact sensitive information (Mendels et al., 2018; Amazon Web Services, Inc., 2025).
@@ 2.2 PII Leakage Attacks
Attacks Toxonomy. According to Lukas et al., 2023 proposed, PII leakage attacks can be categorized into three types based on the adversary’s capabilities: PII extraction, reconstruction, and inference attacks. Extraction attackers have no auxiliary information, while reconstruction attackers have knowledge of masked training texts (e.g., scrubbed datasets) and aim to recover the missing PII. Inference attacks go a step further, where the attacker possesses both masked training texts and PII candidates. The goal of inference is to associate a piece of PII with a given context.
Threat Model. This paper focuses on PII reconstruction attacks because obtaining masked training data is practical in real-world scenarios, and such data allows attackers to exploit contextual dependencies to recover masked PII entities. We assume an adversary with black-box API access to the victim LLM, allowing them to request a continuation or query next-token probabilities to compute cross-entropy loss and PPL. Since API calls entail monetary costs, the adversary should minimize request or query times. Additionally, we assume the adversary knows the reference model, i.e., the base model on which the victim LLM is fine-tuned. It is a reasonable assumption since modern LLM development is usually based on open-source LLMs. And according to popular open-source LLM licenses, such as OpenRAIL-M (BigCode, 2022) and Apache-2.0 Apache (2004), the developer needs to manifest the source when employing a modification of the LLM for commercial use.
@@ 3.2 Candidate Generation
The most widely used previous candidate generation method is TAB (Inan et al., 2021), which generates completions using the true-prefix. TAB behavior like the Tab autocompletion method which automatically fills in table entries, mimicking the behavior of the Tab key. However, its major limitation is that it cannot leverage information from the text following the [MASK]. For instance, in the sentence: “The shipment for [MASK] was delivered to 1234 Elm Street.” TAB fails to utilize the address “1234 Elm Street” to infer the missing PII. Additionally, if other masks appear in the preceding text, TAB requires truncation to prevent the LLM from generating mask tokens. For example, if “shipment” is also masked in the above sentence, TAB can only use “for” as the continuation prefix. Lukas et al., 2023 addressed this issue by filling in previous masks using a Masked Language Model (MLM). However, this approach assumes that the token length of the masked PII is known in advance, which is often impractical in real-world scenarios. Conversely, recollection maximizes the use of both preceding and following context while aligning with the LLM’s training paradigm.
@@ 3.3 Candidate Selection
Theorem 3.1 indicates that our proposed criterion $C$ in Equation 4 can maintain the overlapped correct PII entities predicted by $\mathcal{L}$ and $C_{r}$. We present the detailed proof of Theorem 3.1 in the appendix. When $b\to 0$, $C\to C_{r}$, through Taylor expansion,
@@ 4.1 Experimental Setup
Table 1: Top-1 accuracy of $\mathsf{R.R.}$ and baselines across different models and datasets
@@ 4.4 Ablation Studies
In this section, we conduct ablation studies on two steps of $\mathsf{R.R.}$. For candidate generation and candidate selection, we replace Recollect and $C$ in Equation 4 with alternative methods, respectively.
Ablation of Candidate Generation. We compare Recollect with DirectPrompt and TAB as alternative candidate generation methods. For each approach, we query the victim model 50 times and compare their PII recall. The experimental results are presented in Figure 4. We observe that both Recollect and DirectPrompt showcase an increasing trend in PII recall as the number of query iterations increases. However, TAB reaches a peak with the first few iterations and then stabilizes, indicating that the candidates generated by TAB remain relatively fixed over multiple queries. This also explains why P2P, while doing multiple true-prefix queries, fails to improve prediction accuracy. In contrast, Recollect generally increases PII recall at a faster rate than DirectPrompt and continues to show an upward trend even after 50 iterations, demonstrating the scalability of $\mathsf{R.R.}$.
@@ 4.5 Analysis
We explore how PII entity type and text length affect the reconstruction accuracy in this section. We also analyze the impact of PII occurrence frequency. Due to space limitation, we defer it into Appendix C.
@@
PII Entity Type. We select three PII entity types that appear more than 100 times across all datasets, namely DATA_TIME, PERSON, and LOCATION, to evaluate $\mathsf{R.R.}$’s effectiveness on specific entity types. Figure 5 illustrates the results. Overall, $\mathsf{R.R.}$ demonstrates a strong ability to reconstruct LOCATION entities, while its performance on PERSON entities is weaker. We suspect this is due to the discrepancy in value space; names vary significantly, whereas the number of locations is relatively limited, especially for major cities like New York and Los Angeles, which appear frequently and are easy to guess. Additionally, we observe that DATE_TIME exhibits varying prediction accuracies across different datasets. We suspect this is because DATE_TIME precision differs across datasets. For instance, ECHR often specifies exact dates, while LLM-PC typically records only the year.
@@ 6 Conclusion
In this paper, we propose $\mathsf{R.R.}$ (Recolect and Rank), a novel two-step PII construction attack method. First, we employ a recollection-based method to recover the original text and leverage a NER-model-powered PII identifier to extract potential PII candidates. To evaluate the performance of PII prediction and PII candidate generation, we use top-1 accuracy and PII recall, respectively. Second, we rank the PII candidates by computing their cross-entropy scores under the victim model and a reference model. We further introduce a biased ranking criterion that effectively integrates reference-based and non-reference-based ranking methods. Our experiments across three popular PII datasets and four open-source LLMs demonstrate that the $\mathsf{R.R.}$ significantly improves top-1 accuracy.
Limiations
@@
Lack defense mechanism considerations. Moreover, we didn’t consider the defense mechanisms that may be applied to LLMs, such as differential privacy, adversarial training and other privacy-preserving techniques. These mechanisms are specifically designed to mitigate privacy risks by reducing the amount of sensitive information exposed during model training or inference. While these defenses could reduce the effectiveness of our approach, they may also result in a loss of model utility. Further studies are needed to evaluate the performance of our method on models equipped with such defenses.
@@
$\mathcal{L}(M_{j})\mathcal{L}_{r}(M_{k})$ $<\mathcal{L}(M_{k})\mathcal{L}_{r}(M_{j}).$
Combined with $\mathcal{L}(M_{j})<\mathcal{L}(M_{k})$ and $b>0$, we can get
@@
which completes the proof. $\Box$
@@ B.1 Dataset Details
LLM-PC. This dataset is released by NeuIPS in its 2024 LLM Privacy challenge. LLM-PC consists of two subsets: a development set for assessment and a test set for competition ranking. Each subset contains 1,500 synthetic texts. We use the development set for training victim models and performing PII reconstruction attacks since the test set’s ground truth is not available.

### v2

@@ Abstract
Large Language Models (LLMs) pose significant privacy risks, potentially leaking training data due to implicit memorization. Existing privacy attacks primarily focus on membership inference attacks (MIAs) or data extraction attacks, but reconstructing specific personally identifiable information (PII) in LLMs’ training data remains challenging. In this paper, we propose $\mathsf{R.R.}$ (Recollect and Rank), a novel two-step privacy stealing attack that enables attackers to reconstruct PII entities from scrubbed training data where the PII entities have been masked. In the first stage, we introduce a prompt paradigm named recollection, which instructs the LLM to repeat a masked text but fill in masks. Then we can use PII identifiers to extract recollected PII candidates. In the second stage, we design a new criterion to score each PII candidate and rank them. Motivated by membership inference, we leverage the reference model as a calibration to our criterion. Experiments across three popular PII datasets demonstrate that the $\mathsf{R.R.}$ achieves better PII identification performance than baselines. These results highlight the vulnerability of LLMs to PII leakage even when training data has been scrubbed. We release our code and datasets at GitHub. [^1] https://github.com/meng-wenlong/RR
@@ 1 Introduction
Large Language Models (LLMs) have been widely adopted in various applications, like automated content generation (Girotra et al., 2023; Roziere et al., 2023) to personalized virtual assistants (Nam et al., 2024; Shao et al., 2024). This success is driven by massive training datasets, which raise significant concerns about privacy leakage. Previous research has shown that LLMs can memorize portions of their training data, even before overfitting (Carlini et al., 2019). These training datasets of LLM may contain sensitive personally identifiable information (PII), such as names, phone numbers, and email addresses. Additionally, LLMs are usually fine-tuned with custom datasets to meet application needs, which may introduce further privacy risks.
Figure 1: Comparison of PII prediction accuracies on NeurIPS LLM-PC dataset. $\mathsf{R.R.}$ achieves an improvement of over 100% compared to previous SOTAs.
@@
Recently, Lukas et al., 2023 proposed PII reconstruction attack, which refers to the task where an attacker attempts to recover the masked PII entities in a piece of text. For example, in the masked sentence “The shipment for [MASK] was delivered to 1234 Elm Street,” if the attacker successfully reconstructs [MASK], they can identify the individual residing at that address. This practical attack has attracted great attention in the community. Scrubbed datasets publicly released by LLM developers may serve as potential targets for such attacks. Nevertheless, existing PII reconstruction methods based on prefix continuation (Inan et al., 2021) or perplexity scoring (Lukas et al., 2023), lack contextual dependencies and insights from the original LLM, yielding a reconstruction accuracy below 10%.
Our Contributions. In this paper, we propose $\mathsf{R.R.}$ (Recollect and Rank), a novel and effective PII reconstruction method that assumes black-box access to the victim LLM. $\mathsf{R.R.}$ consists of two stages: (1) recollection, generating PII candidates, and (2) ranking, identifying the candidate most likely to be PII from candidates.
@@
Experimental results of four LLMs show that $\mathsf{R.R.}$ achieves an average top-1 accuracy of 25.73% on the NeurIPS LLM-PC dataset (Li et al., 2024), marking a more than 100% improvement over previous SOTAs, as presented in Figure 1. In summary, our contributions are as follows:
@@ 2.1 LLM Training and PII
LLM Training. We train LLMs using next-token prediction, where the model learns to maximize the probability of correctly predicting the next token given a sequence of preceding tokens. The objective is formulated to minimize cross-entropy loss. Given a sequence of tokens $X=\left(x_{1},x_{2},\ldots,x_{T}\right)$, the cross-entropy loss is defined as:
@@
Personally Identifiable Information (PII). PII refers to any data that can be used to uniquely identify an individual, such as names, addresses, phone numbers, or social security numbers. Since LLMs are often trained on vast real-world texts, they inevitably ingest and memorize a significant amount of PII. To detect PII in natural language text, Named Entity Recognition (NER) models are commonly employed. NER is a token classification task, where each token in a sequence is assigned a label corresponding to a specific entity type. Several industry-grade PII detection services utilize NER-based models to identify and redact sensitive information (Mendels et al., 2018; Amazon Web Services, Inc., 2025).
@@ 2.2 PII Leakage Attacks
Attacks Taxonomy. According to Lukas et al., 2023 proposed, PII leakage attacks can be categorized into three types based on the adversary’s capabilities: PII extraction, reconstruction, and inference attacks. Extraction attackers have no auxiliary information, while reconstruction attackers have knowledge of masked training texts (e.g., scrubbed datasets) and aim to recover the missing PII. Inference attacks go a step further, where the attacker possesses both masked training texts and PII candidates. The goal of inference is to associate a piece of PII with a given context.
Threat Model. This paper focuses on PII reconstruction attacks because obtaining masked training data is practical in real-world scenarios. We assume an adversary with black-box API access to the victim LLM, allowing them to request a continuation or query next-token probabilities to compute cross-entropy loss and PPL. Since API calls entail monetary costs, the adversary should minimize request or query times. Experimental results (Section 4.4) prove that our $\mathsf{R.R.}$ requires fewer queries to achieve the same PII recall in candidate generation. In candidate selection, we neglect the content preceding PII when computing loss, reducing the query count. Additionally, we assume the adversary knows the reference model, i.e., the base model on which the victim LLM is fine-tuned. It is a reasonable assumption since modern LLM development is usually based on open-source LLMs. And according to popular open-source LLM licenses, such as OpenRAIL-M (BigCode, 2022) and Apache-2.0 Apache (2004), the developer needs to manifest the source when employing a modification of the LLM for commercial use. Even if the attacker uses an incorrect reference model, we show that it causes a small degradation in reconstruction accuracy (Appendix C).
@@ 3.2 Candidate Generation
The most widely used previous candidate generation method is TAB (Inan et al., 2021), which generates completions using the true prefix. TAB behaves like the Tab autocompletion method which automatically fills in table entries, mimicking the behavior of the Tab key. However, its major limitation is that it cannot leverage information from the text following the [MASK]. For instance, in the sentence: “The shipment for [MASK] was delivered to 1234 Elm Street.” TAB fails to utilize the address “1234 Elm Street” to infer the missing PII. Additionally, if other masks appear in the preceding text, TAB requires truncation to prevent the LLM from generating mask tokens. For example, if “shipment” is also masked in the above sentence, TAB can only use “for” as the continuation prefix. Lukas et al., 2023 addressed this issue by filling in previous masks using a Masked Language Model (MLM). However, this approach assumes that the token length of the masked PII is known in advance, which is often impractical in real-world scenarios. Conversely, recollection maximizes the use of both preceding and following context while aligning with the LLMs’ training paradigm.
@@ 3.3 Candidate Selection
Theorem 3.1 indicates that our proposed criterion $C$ in Equation 4 can maintain the overlapped correct PII entities predicted by $\mathcal{L}$ and $C_{r}$. We present the detailed proof of Theorem 3.1 in the appendix. When $b\to 0$, $C\to C_{r}$. Through Taylor’s expansion,
@@ 4.1 Experimental Setup
Table 1: Top-1 accuracy of $\mathsf{R.R.}$ and baselines across different models and datasets.
@@ 4.3 Effectiveness of Biased Criterion
(a) Comparison of PII recalls on GPT-4o-mini.
(b) Prompt refusal rates on proprietary models.
Figure 5: Experiments on real-world models.
@@ 4.4 Ablation Studies
In this section, we conduct ablation studies on two steps of $\mathsf{R.R.}$. For candidate generation and candidate selection, we replace Recollect and $C$ in Equation 4 with alternative methods, respectively. We also compare hybrid attacks that combine different candidate generation and candidate selection methods. Due to the space limitation, we defer hybrid attack results to Appendix D.
Ablation of Candidate Generation. We compare Recollect with DirectPrompt and TAB as alternative candidate generation methods. For each approach, we query the victim model 50 times and compare their PII recall. The experimental results are presented in Figure 4. We observe that both Recollect and DirectPrompt showcase an increasing trend in PII recall as the number of query iterations increases. However, TAB reaches a peak with the first few iterations and then stabilizes, indicating that the candidates generated by TAB remain relatively fixed over multiple queries. This also explains why P2P, while doing multiple true-prefix queries, fails to improve prediction accuracy. In contrast, Recollect generally increases PII recall at a faster rate than DirectPrompt and continues to show an upward trend even after 50 iterations, demonstrating the scalability of $\mathsf{R.R.}$. We also conduct an experiment on GPT-4o-mini using OpenAI Fine-tuning API to demonstrate that our recollection method remains powerful to commercial models. We randomly select 50 texts from LLM-PC, containing 218 PII entities and fine-tube GPT-4o-mini for 5 epochs. The PII recall versus the number of query iterations is shown in 5(a). Recollect achieves the highest PII recall when giving the same query budget.
Another concern is that victim models may refuse to output or provide perturbed results when detecting privacy-related queries. To address this concern, we conduct an experiment using 200 prompts for each candidate generation method to investigate whether current proprietary models would refuse them. The refusal rate results are shown in 5(b). We observe that DirectPrompt occasionally triggers malicious behavior detection, causing the model to refuse to respond. TAB sometimes results in refusals from GPT-4o and GPT-4o-mini due to too short context provided for continuation. In contrast, none of the above four proprietary models refuses Recollect. This may be because Recollect disguises itself by presenting the task of filling the missing parts.
Table 3: comparison of model performance and PII prediction accuracy for models trained with DP-SGD under varying privacy budget ($\epsilon$).
Llama32-3B Phi-3.5-mini
$\boldsymbol{\epsilon}$ 4 8 16 32 $\infty$ 4 8 16 32 $\infty$
Train Loss 2.50 2.48 2.47 2.46 1.79 2.15 2.12 2.13 2.03 1.81
ACC (DirectPrompt) 0.22% 0.39% 0.57% 0.44% 3.56% 0.44% 0.48% 0.84% 0.84% 2.11%
ACC (TAB) 0.53% 0.35% 0.53% 0.62% 7.20% 0.40% 0.31% 0.44% 0.53% 4.09%
ACC (P2P) 0.57% 0.62% 0.62% 0.75% 6.91% 0.48% 0.53% 0.58% 0.62% 4.28%
ACC (R.R.) 3.60% 3.74% 3.87% 4.18% 14.79% 2.95% 3.21% 3.64% 3.91% 11.10%
@@
4.5 Defending against $\mathsf{R.R.}$
To mitigate the risk of PII leakage from LLMs, developers may employ privacy-preserving techniques. The most common and widely adopted approach is Differential Privacy (DP) (Dwork, 2006), which provides rigorous mathematical guarantees that an individual’s data cannot be inferred through differential attacks. DP-SGD (Abadi et al., 2016) is the most prominent method for training deep learning models to satisfy DP. Given a privacy budget ($\epsilon$), DP-SGD operates by clipping per-sample gradients and adding scaled noise during each update. In this work, we fine-tune victim models on the ECHR dataset using DP-SGD implemented by Microsoft’s DP-Transformers (Wutschitz et al., 2022). Since DP-SGD incurs extra memory overhead and is hard to integrate with distributed learning, we only conduct DP-SGD experiments on 3B models. The results are shown in Table 3.
We find that although DP-SGD effectively reduces reconstruction accuracy, it also leads to a significant increase in loss, jeopardizing model performance. Our baselines become nearly unusable on DP fine-tuned models, whereas $\mathsf{R.R.}$ remains about 1/3 accuracy. The results suggest that while DP-SGD is a powerful privacy-preserving method, it cannot completely defend against our $\mathsf{R.R.}$. An alternative defense is to prevent the model from learning PII tokens during training. For instance, Lin et al., 2024 proposed Selective Language Modeling (SLM), which selectively trains on useful tokens. We hypothesize that most PII tokens contribute little to the general language modeling capability of LLMs. Thus, selectively excluding them could enhance privacy with minimal impact on utility. Another line of defense is post-hoc unlearning (Yao et al., 2024; Liu et al., 2024). The developer can remove specific PII from LLMs after training using unlearning techniques.
@@ 4.6 Analysis
We explore how PII entity type and text length affect the reconstruction accuracy in this section. We also analyze the impact of PII occurrence frequency. Due to space limitation, we defer it into Appendix E.
@@
PII Entity Type. We select three PII entity types that appear more than 100 times across all datasets, namely DATE_TIME, PERSON, and LOCATION, to evaluate $\mathsf{R.R.}$’s effectiveness on specific entity types. Figure 6 illustrates the results. Overall, $\mathsf{R.R.}$ demonstrates a strong ability to reconstruct LOCATION entities, while its performance on PERSON entities is weaker. We suspect this is due to the discrepancy in value space; names vary significantly, whereas the number of locations is relatively limited, especially for major cities like New York and Los Angeles, which appear frequently and are easy to guess. Additionally, we observe that DATE_TIME exhibits varying prediction accuracies across different datasets. We suspect this is because DATE_TIME precision differs across datasets. For instance, ECHR often specifies exact dates, while LLM-PC typically records only the year.
@@ 6 Conclusion
In this paper, we propose $\mathsf{R.R.}$ (Recollect and Rank), a novel two-step PII reconstruction attack. First, we employ a recollection-based method to recover the original text and leverage a NER-model-powered PII identifier to extract potential PII candidates. Then, we rank the PII candidates by computing their cross-entropy scores under the victim model and a reference model. We further introduce a biased ranking criterion that effectively integrates reference-based and non-reference-based ranking methods. Our experiments across three popular PII datasets and four open-source LLMs demonstrate that the $\mathsf{R.R.}$ significantly improves top-1 accuracy.
Acknowledgements
This research received partial support from the National Natural Science Foundation of China under Grant No. 62302441. This work was also supported by the Key Research and Development Program Project of Ningbo Grant No. 2025Z029. Infrastructure support was provided by the Information Technology Center of Zhejiang University and the Supercomputing Center of Hangzhou City University.
Limitations
@@ Acknowledgements
Robustness against Defense Mechanisms. In real-world applications, developers often employ privacy-preserving techniques, such as differential privacy, to effectively mitigate potential risks of PII leakage. As demonstrated in  Section 4.5, DP-SGD significantly reduces the PII prediction accuracy of our proposed method to one-third. However, despite this degradation, our proposed method still outperforms other baseline approaches. Furthermore, such defense mechanisms not only incur higher computational costs, but also reduce the utility of LLMs.
@@
$\mathcal{L}(M_{j})\mathcal{L}_{r}(M_{k})$ $<\mathcal{L}(M_{k})\mathcal{L}_{r}(M_{j}).$ (8)
Because $b$ is a positive bias, we can get
$b\mathcal{L}\left(M_{j}\right)<b\mathcal{L}\left(M_{k}\right).$ (9)
Combing Appendix A and Equation 9 we obtain
@@
Based on our assumption that $k$ can be any other candidate index that is unequal to $j$, $M_{j}$ will also be ranked as the first by $\frac{\mathcal{L}-\left(\mathcal{L}_{r}+b\right)}{\mathcal{L}_{r}+b}$. $\Box$
@@ B.1 Dataset Details
LLM-PC. This dataset is released by NeurIPS in its 2024 LLM Privacy challenge. LLM-PC consists of two subsets: a development set for assessment and a test set for competition ranking. Each subset contains 1,500 synthetic texts. We use the development set for training victim models and performing PII reconstruction attacks since the test set’s ground truth is not available.